# 4D Millimeter-Wave Radar in Autonomous Driving: A Survey

### v3

@@ Abstract
The 4D millimeter-wave (mmWave) radar, proficient in measuring the range, azimuth, elevation, and velocity of targets, has attracted considerable interest within the autonomous driving community. This is attributed to its robustness in extreme environments and the velocity and elevation measurement capabilities. However, despite the rapid advancement in research related to its sensing theory and application, there is a conspicuous absence of comprehensive surveys on the subject of 4D mmWave radar. In an effort to bridge this gap and stimulate future research, this paper presents an exhaustive survey on the utilization of 4D mmWave radar in autonomous driving. Initially, the paper provides reviews on the theoretical background and progress of 4D mmWave radars, encompassing aspects such as the signal processing workflow, resolution improvement approaches, and extrinsic calibration process. Learning-based radar data quality improvement methods are present following. Then, this paper introduces relevant datasets and application algorithms in autonomous driving perception, localization and mapping tasks. Finally, this paper concludes by forecasting future trends in the realm of 4D mmWave radar in autonomous driving. To the best of our knowledge, this is the first survey specifically dedicated to the 4D mmWave radar in autonomous driving.
@@ I Introduction
Autonomous driving technology, which aspires to provide safe, convenient, and comfortable transportation experiences, is advancing at a remarkable pace. To realize high-level autonomous driving, the capabilities of environment perception, localization and mapping are indispensable. Consequently, the sensors deployed on autonomous vehicles, such as cameras, LiDARs, and radars, along with their application algorithms, are garnering increasing research interest.
@@
Refer to caption (a)
Refer to caption (b)
Figure 1: The Oculii Eagle 4D mmWave radar (a) and its point cloud comparing with the Ouster 128-channel LiDAR (b) [2, 3]
@@
The recent advancement of multiple-input multiple-output (MIMO) antenna technology has catalyzed a significant enhancement in elevational resolution, leading to the emergence of 4D mmWave radars. As the name suggests, 4D mmWave radars are capable of measuring four distinct types of target information: range, azimuth, elevation, and velocity. In addition to the augmented elevational resolution, 4D mmWave radars still preserve the salient advantages of their 3D predecessors. Enterprises that involve in the 4D mmWave radar industry ranging from conventional suppliers like Bosch, Continental, and ZF, to a host of burgeoning tech companies such as Arbe, Huawei, and Oculii. An illustrative example of this technology is the Oculii Eagle 4D mmWave radar, which, when juxtaposed with the Ouster 128-channel LiDAR, demonstrates its capabilities such as long detection range through the point cloud representation as depicted in Fig. 1.
The 4D mmWave radar not only represents an improved version of mmWave radars but also ushers in a multitude of significant research avenues. The raw data volume generated by 4D mmWave radars substantially exceeds that of their traditional counterparts, thereby presenting formidable challenges in signal processing and data generation, particularly in the presence of clutter and noise. The sparsity and noise inherent in 4D mmWave radar point clouds, generated in the existing signal processing workflow are markedly more severe than those in LiDAR point clouds. This necessitates the careful design of perception, localization and mapping algorithms that duly accommodate the intrinsic characteristics of the 4D mmWave radar.
@@
Numerous surveys have been conducted on the theory and application of mmWave radars. In recent years, Bilik et al. [4] have reviewed the challenges faced by mmWave radars in autonomous driving and its future trends. Venon et al. [5] have provided a comprehensive summary of the theory and existing perception algorithms of mmWave radar in autonomous driving, while Harlow et al. [6] have concentrated on mmWave radar applications in robotics for their survey. It is evident that most reviews are centered on 3D mmWave radars.
Despite the transformative emergence of 4D mmWave radars and associated algorithms, there have been few specialized surveys. To bridge this gap, this paper presents a thorough review of 4D mmWave radars in autonomous driving. The principal contributions of this work can be summarized as follows:
  • To the best of our knowledge, this is the first survey concentrating on 4D mmWave radars within the context of autonomous driving.
  • Acknowledging the distinctiveness of 4D mmWave radars, this survey delineates its theoretical background, signal processing workflow, and discusses newly emerged learning-based radar data quality improvement methods.
  • This paper delivers an extensive survey of 4D mmWave radar application algorithms in autonomous driving. It systematically presents related datasets, as well as research on perception, localization and mapping algorithms of 4D mmWave radars, all of which are chronologically depicted on a timeline in Fig. 2.
@@
Figure 2: The timeline of 4D mmWave radar datasets, learning-based radar data quality improvement methods, and perception, localization and mapping algorithms
Figure 3: The categories of newly evolved 4D mmWave radar technologies in autonomous driving.
@@
The remainder of this paper is organized as follows: Section II introduces the foundational theory of 4D mmWave radars, including the signal processing workflow, data formats, and the methods for improving resolution and calibration. Following Sections illustrate newly evolved 4D mmWave radar technologies in autonomous driving, as categorised in Fig. 3. Section III lists available 4D mmWave radar datasets for researchers’ convenience. Section IV summarizes some learning-based methods for radar data quality improvement. Section V reviews 4D mmWave radar perception applications, categorized into different input formats. Moreover, 4D mmWave radar applications in localization and mapping are presented in Section VI. Section VII discusses future trends of 4D mmWave radar in autonomous driving, and Section VIII draws the conclusion.
@@ II-C Extrinsic Calibration
Following traditional calibration methods of 3D mmWave radars, corner reflectors are commonly employed to improve calibration accuracy. By carefully placing several corner reflectors and analyzing the sensing results of the 4D mmWave radar in conjunction with LiDAR and camera data, the extrinsic parameters can be calibrated [17]. In a departure from the conventional approach of calibrating each sensor pair sequentially, Domhof et al. calibrate all sensors simultaneously with respect to the body of mobile robot, achieving a median rotation error of a mere 0.02 ${}^{\circ}$ [18]. By leveraging the Random Sample and Consensus (RANSAC) and Levenberg-Marquardt non-linear optimization, [19] accomplishes radar-camera calibration with only one single corner reflector, obviating the requirement of a specially designed calibration environment.
However, the practicability of corner reflectors in real-world scenarios is limited. Recent research has proposed calibration methods for 4D mmWave radars that eschew the need for specially placed corner reflectors, instead utilizing radar motion measurement to conduct online calibration for radars [20] or radar-camera pairs [21]. While these methods offer convenience, their efficacy under extreme weather conditions remains to be validated.
@@
III Datasets
Public datasets play an indispensable role in the advancement of 4D mmWave radar-based algorithms, as they furnish essential platforms for the development, benchmarking, and comparative analysis of diverse algorithms, thereby stimulating research in the field. Current available datasets with 4D mmWave radars are introduced in this section, which are summarized in Table I.
TABLE I: 4D mmWave radar datasets
[…]
Azi. Ele. Range(m) Velo.(m/s)
[…]
SCORP [26] ${}^{\circ}$ ${}^{\circ}$ 12 0.33 3.9K 3.9K ADC, RT, RPC RC $\times$ $\times$
Coloradar [10] ${}^{\circ}$ ${}^{\circ}$ 0.12 0.25 108K 0K ADC, RT, RPC RLI $\times$ $\times$
Radatron [27] ${}^{\circ}$ ${}^{\circ}$ 0.05 N/M 152K 16K ADC RC 2D $\times$ $\times$
RADIal [28] ${}^{\circ}$ ${}^{\circ}$ 0.2 0.1 25K 8.3K ADC, RT, RPC RCL 2D $\times$
RPDNet [29] ${}^{\circ}$ ${}^{\circ}$ 0.03 0.03 28K 0K RT RCLI 3D $\times$
VoD [30] ${}^{\circ}$ ${}^{\circ}$ 0.2 0.1 8.7K 8.7K RPC RCI 3D
TJ4DRadSet [17] ${}^{\circ}$ ${}^{\circ}$ 0.86 N/M 40K 7.8K RPC RCL 3D
K-Radar [31] ${}^{\circ}$ ${}^{\circ}$ 0.46 0.06 35K 35K RT, RPC RCLI 3D
MSC-RAD4R [3] ${}^{\circ}$ ${}^{\circ}$ 0.86 0.27 90K 0K RPC RCLI $\times$ $\times$
NTU4DRadarLM [32] ${}^{\circ}$ ${}^{\circ}$ 0.86 N/M 61K 0K RPC RCLIT $\times$ $\times$
Dual Radar [33] ${}^{\circ}$ ${}^{\circ}$ 0.22 N/M 50K 10K RPC RCL 3D $\times$
  • ${}^{1}$: raw radar data after Analog-to-Digital Converter; RT: radar tensor representations like RA,RD,RAD or 4DRT; RPC: radar point cloud;
  • ${}^{2}$: Radar, C:Camera, L:LiDAR, I:IMU(Inertial Measurement Unit), T:Thermal Camera
  • ${}^{3}$/M: Not Mentioned.
We have curated published datasets according to the specific tasks they are tailored for, aiming to provide a comprehensive and systematically organized overview of datasets to aid the research and development of 4D mmWave radar-based algorithms.
III-A Datasets for Perception
Datasets designed for 4D mmWave radar perception typically include 3D (or 2D) bounding boxes for object detection tasks, and tracking ID for object tracking tasks.
Public available 4D mmWave radar dataset are relatively rare. Astyx [25] represents the first dataset of its kind. The data consists of 500 synchronized frames (radar, LiDAR, camera) encompassing approximately 3,000 annotated 3D object annotations. As a pioneer dataset in the realm of 4D mmWave radar, the volume of data in Astyx is relatively limited. The Radatron dataset[27] enhances the data resolution through a cascaded MIMO radar and provides radar heatmaps with 2D bounding boxes labeled by camera ground truth. However, as Radatron was collected under clear weather conditions, it is not much suitable for demonstrating radar performance under severe weather conditions.
In order to facilitate researchers in handling radar data in a more fundamental manner, the RADIal dataset [28] records sensor signals in their raw format. This raw signal corresponds to the ADC output. This ADC data serves as the foundation for generating various conventional radar representations, such as radar tensors and point clouds. Given that the raw ADC data is not interpretable by human eyes, the annotations in the RADIal dataset are presented as 2D bounding boxes in the image plane.
The VoD dataset [30] is a multi-sensor automotive dataset for multi-class 3D object detection, comprising calibrated and synchronized LiDAR, camera, and radar data. It contains 8693 frames of data captured in complex urban traffic scenarios, and includes 123106 3D bounding box annotations of both stationary and dynamic objects and tracking IDs for each annotated object. Road user detection is the primary concern of the VoD dataset, which therefore differentiates between pedestrian, cyclist, and car labels.
In a similar vein, the Tj4DRadSet dataset [17] comprises 44 consecutive sequences, totaling 7757 synchronized frames, well-labeled with 3D bounding boxes and trajectory IDs. Notably, the TJ4DRadSet dataset offers occlusion and truncation indicators in each object to distinguish between different levels of detection difficulty. This dataset encompasses eight categories: car, bus, truck, construction vehicle, pedestrian, motorcyclist, bicyclist, and tricyclist. Unlike the VoD dataset, Tj4DRadSet is characterized by its inclusion of a broader and more challenging array of driving scenario clips, such as urban roads, highways, and industrial parks.
To the best of our knowledge, K-Radar [31] currently contains most diverse scenarios in 4D mmWave radar datasets as it collects 35k frame under a variety of weather conditions, including sunny, foggy, rainy, and snowy. K-Radar not only provides 4D mmWave radar data but also includes high-resolution LiDAR point clouds, surround RGB imagery from four stereo cameras, and RTK-GPS and IMU data from the ego-vehicle. Fig. 6 illustrates the comparison of different modality sensors across different weather conditions. It is worth mentioning that K-radar is currently the only dataset that provides 4D mmWave radar tensors. In order to facilitate experiments with various neural network architectures, K-radar also provides a visual programming tool to modularize the neural network training code.
However, the 4D mmWave radar models employed in the aforementioned datasets are inconsistent, making it challenging for researchers to analyze and compare the performance of different 4D mmWave radar data. The recently unveiled Dual Radar dataset[33], as illustrated in Figure 7, encompasses two distinct types of mmWave radars, the Arbe Phoenix and the ARS548 RDI radar. The collection of 4D mmWave radar point clouds from these two radar types within identical scenes enables an investigation into the impact of different sparsity levels in point clouds on object detection performance. This exploration is poised to lay the groundwork for furthering research on 4D mmWave radar technology in the autonomous vehicle sector.
Figure 7: The data visualization of Dual Radar dataset. [33]
III-B Datasets for Localization and Mapping
4D mmWave radar is capable of generating dense 3D point clouds that incorporates elevation information, thereby facilitating robust localization even under challenging environmental conditions. In recent years, there has been a growing recognition of the critical significance of 4D mmWave radar in localization and mapping, leading to the public release of several 4D mmWave radar datasets specifically designed for these tasks. Odometry information is indispensable for localization, mapping, or Simultaneous Localization and Mapping (SLAM). Therefore, besides datasets specialized for localization and mapping, datasets introduced above can also be employed, provided they include odometry information. 4D mmWave radar technology is capable of generating a dense 3D point cloud that incorporates elevation information, thereby facilitating robust localization even under challenging environmental conditions. In recent years, the critical importance of 4D mmWave radar in localization and mapping has gained increasing recognition, leading to the public release of several 4D mmWave radar datasets specifically designed for these tasks. Odometry information is crucial for localization, mapping, or Simultaneous Localization and Mapping (SLAM). In addition to datasets specifically designed for these tasks, the datasets mentioned above can also be utilized for localization and mapping, provided they include odometry information.
ColoRadar [10] is a dataset dedicated to research on localization and mapping, comprising approximately 2 hours of data from radar, LiDAR, and the 6-DOF pose ground truth. It provides radar data of three processing levels: raw ADC data, 3D range-azimuth-elevation tensors derived by compressing the Doppler dimension of 4D mmWave radar tensors, and radar point clouds. This dataset collects data from a variety of unique indoor and outdoor environments, thus providing a diverse spectrum of sensor data. In the Coloradar dataset, the radar is mounted on a robot with a relatively restrained mobility profile without GNSS integration, thereby limiting access to ground truth information regarding the robot’s position.
The MSC-RAD4R dataset [3] is a ROS-based 4D mmWave radar dataset for SLAM. It records data under a wide range of weather and environmental conditions, with the same route yielding data from both clear and snowy weather conditions for comparison. Additionally, MSC-RAD4R introduces artificially generated smoke environment data generated by a smoke machine, further emphasizing the robust capabilities of 4D mmWave radars.
The NTU4DRadLM dataset [32] is a recent contribution to this field. Distinguished from its predecessors, NTU4DRadLM delivers an extensive array of localization-relevant sensor data, including 4D mmWave radar, LiDAR, camera, IMU, GPS, and even a thermal camera. Curated explicitly for SLAM tasks, it provides fine-tuned ground truth odometry and supports closed-loop trajectory configurations to enable potential loop detection and back-end optimization. Accommodating varying sensor platform velocities, NTU4DRadLM’s data is captured using both robotic and vehicular platforms. Furthermore, it encompasses a wide range of road conditions, encompassing structured, semi-structured, and unstructured roads, spanning both small-scale environments (e.g., gardens) and large-scale urban settings.
In addition, several datasets enumerated in Table I provide radar data with relatively low angular resolution [26, 29]. These datasets, which are not detailed within the scope of this paper, provide alternative radar data characteristics that may be leveraged in different research contexts.
III-C Challenge
Taking into account the entirety of the datasets delineated afore, it becomes apparent that there is a lack of labeled radar ADC data required for deep radar detection. One possibility to overcome this deficiency is the employment of syntactic data generation methods that relies on some sensor model. Unfortunately, radar modeling is extremely challenging and computationally intensive due to the multiple effects that occur in acquisition, such as multi-path reflections, signal interference, reflective surface interactions, discrete resolution cells, and signal attenuation.
Moreover, current 4D mmWave radar datasets do not rival the voluminous scale of other autonomous driving datasets such as nuScenes[34] and ONCE[35]. The substantial size of datasets is imperative for the evaluation of algorithmic generalizability and for facilitating comparative analysis with other sensor modalities in the context of 4D mmWave radar research.
IV Learning-based Radar Data Enhancement
Despite the extensive development of radar imaging technology over the years, traditional handcrafted methods such as CFAR continue to manifest inherent limitations. These methods, while foundational, often fail to account for the diverse and complex nature of real-world targets, which can vary significantly in shape and extend across multiple-resolution cells. Such discrepancies can induce masking effects within CFAR-type algorithms, leading to a diminution in the resolution of point clouds and consequent information loss.
In response to the limitations of conventional methods, this section will pivot to exploring learning-based radar data quality improvement techniques. The advent of deep learning, particularly within the autonomous driving domain, offers promising avenues for addressing the deficiencies above. By leveraging the capabilities of deep learning, it is possible to develop more adaptive and robust algorithms that can improve the fidelity of radar imaging. Additionally, we will discuss the challenges we believe still exist in this area, thereby inspiring the research community.

### v4

@@ Abstract
The 4D millimeter-wave (mmWave) radar, proficient in measuring the range, azimuth, elevation, and velocity of targets, has attracted considerable interest within the autonomous driving community. This is attributed to its robustness in extreme environments and the velocity and elevation measurement capabilities. However, despite the rapid advancement in research related to its sensing theory and application, there is a conspicuous absence of comprehensive surveys on the subject of 4D mmWave radars. In an effort to bridge this gap and stimulate future research, this paper presents an exhaustive survey on the utilization of 4D mmWave radars in autonomous driving. Initially, the paper provides reviews on the theoretical background and progress of 4D mmWave radars, encompassing aspects such as the signal processing workflow, resolution improvement approaches, and extrinsic calibration process. Learning-based radar data quality improvement methods are present following. Then, this paper introduces relevant datasets and application algorithms in autonomous driving perception and localization tasks. Finally, this paper concludes by forecasting future trends in the realm of the 4D mmWave radar in autonomous driving. To the best of our knowledge, this is the first survey specifically dedicated to the 4D mmWave radar in autonomous driving.
Index Terms:
4D millimeter-wave radar, Autonomous driving, Perception, SLAM, Dataset
@@ I Introduction
Autonomous driving technology, which aspires to provide safe, convenient, and comfortable transportation experiences, is advancing at a remarkable pace. To realize high-level autonomous driving, the capabilities of environment perception and localization are indispensable. Consequently, the sensors deployed on autonomous vehicles, such as cameras, LiDARs, and radars, along with their application algorithms, are garnering increasing research interest.
@@
Figure 1: The main pipeline of this survey.
@@
The recent advancement of multiple-input multiple-output (MIMO) antenna technology has catalyzed a significant enhancement in elevational resolution, leading to the emergence of 4D mmWave radars. As the name suggests, 4D mmWave radars are capable of measuring four distinct types of target information: range, azimuth, elevation, and velocity. In addition to the augmented elevational resolution, 4D mmWave radars still preserve the salient advantages of their 3D predecessors. The comparison among autonomous driving sensors, including 4D mmWave radar, 3D mmWave radar, LiDAR, RGB camera and thermal camera is shown in Table. I. The 4D mmWave radar distinctly holds advantages in velocity measurement, detection range, all-environment robustness and low cost. Enterprises that involve in the 4D mmWave radar industry ranging from conventional suppliers like Bosch, Continental, and ZF, to a host of burgeoning tech companies such as Arbe, Huawei, and Oculii. An illustrative example of this technology is the Oculii Eagle 4D mmWave radar, which, when compared with the Ouster 128-channel LiDAR, demonstrates its capabilities such as long detection range through the point cloud representation as depicted in Fig. 2.
Figure 2: The Continental ARS548RDI 4D mmWave radar (a), Oculii Eagle 4D mmWave radar (b) and the point cloud of Oculii Eagle comparing with the Ouster 128-channel LiDAR (c). [2]
TABLE I: Comparison among autonomous driving sensors and data formats
Features 4D mmWave Radar 3D mmWave Radar LiDAR RGB Camera Thermal Camera
Range Resolution High High Very High Low Low
Azimuth Resolution High Moderate Very High Moderate Moderate
Elevation Resolution High Unmeasurable Very High Moderate Moderate
Velocity Resolution High High Unmeasurable Unmeasurable Unmeasurable
Detection Range High High Moderate Low Moderate
Surface Measurement Texture Texture No Color Thermal Signature
Lighting Robustness High High High Low High
Weather Robustness High High Low Low High
Cost Moderate Low High Moderate High
However, as a newly developed sensor, the 4D mmWave radar also presents some challenges due to its inherent characteristics. On the one hand, the raw data volume generated by the 4D mmWave radar substantially exceeds that of its traditional counterpart, thereby presenting formidable problems in signal processing and data generation. on the other hand, the sparsity and noise natural in 4D mmWave radar point clouds, generated in the existing signal processing workflow are notably more severe than those in LiDAR point clouds. Aiming at solving these issues, as well as utilizing 4D mmWave radar features such as Doppler and elevation measurement, a great number of researchers have engaged in studies within the fields of 4D mmWave radar-based data generation[3, 4], perception[5, 6] and SLAM(Simultaneous Localization and Mapping)[7, 8].
@@
In recent years, numerous surveys have been conducted on the theory and application of mmWave radars[9, 10, 11, 12, 13, 14], but most of them are centered on 3D mmWave radars. Bilik et al. [9] have reviewed the challenges faced by mmWave radars in autonomous driving and its future trends. Venon et al. [10] have provided a comprehensive summary of the theory and existing perception algorithms of mmWave radar in autonomous driving, while Harlow et al. [11] have concentrated on mmWave radar applications in robotics for their survey.
@@
Figure 3: The timeline of 4D mmWave radar-related works, including learning-based radar data generation methods, perception and SLAM algorithms, and datasets
Despite the transformative emergence of 4D mmWave radars and associated algorithms, there have been few specialized surveys. Liu et al. [15] compare different pipelines of 4D mmWave radar-based object tracking algorithms. Fan et al. [14] summarize perception and SLAM applications for 4D mmWave radar, but ignore the quite important radar data generation studies, and the logical framework within the applications are not systematically outlined. To bridge this gap, this paper presents a thorough review of 4D mmWave radars in autonomous driving. The principal contributions of this work can be summarized as follows:
  • To the best of our knowledge, this is the first publicly available survey focuses on 4D mmWave radars within the context of autonomous driving.
  • Acknowledging the distinctiveness of 4D mmWave radars, this survey outlines its theoretical background, draws a detailed signal processing workflow figure, as the foundation of its application.
  • Given the sparsity and noise of existing 4D mmWave radar point cloud, this paper discusses newly emerged learning-based radar data generation methods that can enhance data quality.
  • This paper delivers an extensive survey of 4D mmWave radar application algorithms in autonomous driving. It systematically presents research on perception and SLAM algorithms of 4D mmWave radars, as well as related datasets, and categorizes them on a timeline in Fig. 3.
  • By thoroughly tracing relevant research, this paper presents classification framework diagrams for 4D mmWave radar data generation, perception, and SLAM applications. Existing challenges and in-depth future outlook are also illustrated.
@@
The remainder of this paper is organized as shown in Fig. 1: Section II introduces the foundational theory of 4D mmWave radars, including the signal processing workflow, methods for improving resolution and extrinsic calibration. Section III summarizes some learning-based methods for radar data generation. Section IV reviews 4D mmWave radar perception applications, categorized into different input formats. 4D mmWave radar applications in SLAM are presented in Section V. Moreover, Section VI lists available 4D mmWave radar datasets for researchers’ convenience. Section VII discusses future trends of 4D mmWave radar in autonomous driving, and Section VIII draws the conclusion.
@@ II-C Extrinsic Calibration
Following traditional calibration methods of 3D mmWave radars, corner reflectors are commonly employed to improve calibration accuracy. By carefully placing several corner reflectors and analyzing the sensing results of the 4D mmWave radar in conjunction with LiDAR and camera data, the extrinsic parameters can be calibrated [25]. In a departure from the conventional approach of calibrating each sensor pair sequentially, Domhof et al. calibrate all sensors simultaneously with respect to the body of mobile robot, achieving a median rotation error of a mere 0.02 [26]. By leveraging the Random Sample and Consensus (RANSAC) and Levenberg-Marquardt non-linear optimization, [27] accomplishes radar-camera calibration with only one single corner reflector, obviating the requirement of a specially designed calibration environment.
However, the practicability of corner reflectors in real-world scenarios is limited. Recent research has proposed calibration methods for 4D mmWave radars that avoid the need for specially placed corner reflectors, instead utilizing radar motion measurement to conduct online calibration for radars [28] or radar-camera pairs [29]. While these methods offer convenience, their efficacy under extreme weather conditions remains to be validated.
@@
III Learning-based Radar Data Generation
As discussed in Section II, the initial data from 4D mmWave radars comprise spectral signals heavily masked by noise, necessitating filtering algorithms like CFAR to generate usable point clouds. However, such traditional handcrafted methods have inherent limitations. They often struggle to adapt to the complexity of real-world targets, which can vary significantly in shape and extend across multiple resolution cells. This mismatch can induce masking effects within CFAR-type algorithms, consequently reducing the resolution of point clouds and resulting in significant information loss.
(a) LiDAR point cloud
(b) RGB image
(c) OS-CFAR point cloud
(d) NN-Generated point cloud
Figure 6: Object detection results from LiDAR point clouds and radar point clouds generated by the OS-CFAR detector and learning-based detector[4].
To overcome those limitations, this section introduces learning-based techniques for radar data generation. By leveraging the capabilities of deep learning, it is possible to develop more adaptive and robust algorithms that can improve the fidelity of radar imaging. As illustrated in Fig. 6, radar point clouds generated using a learning-based detector are denser and contain less noise compared to those produced by the traditional OS-CFAR detector.
In the current landscape of mmWave radar technology, two primary learning-based pipelines have been developed: "Reconstructor" and "Detector," as shown in Fig. 7. Reconstructor techniques focus on refining radar point clouds by enhancing their density and resolution. In contrast, Detector methods bypass the preliminary CFAR filtering stage and process radar frequency data directly, thus preventing the information loss typically associated with traditional filtering methods. Subsequent sections will provide a detailed comparison of these approaches and discuss the persistent challenges in this area.